APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Monday February 9, 2009 (MJD 54872) ---=== OBSERVING TEAM ===--- Howard Brewington Audrey Simmons Steph Snedden (afternoon, support) MARVELS Team - Nathan De Lee and Liming Dou ---=== OBSERVING PLAN ===--- 9 Feb 2009 54872.0 0.95 MARVELS 11.3 hours Flexible Engineering ---=== OBSERVING SUMMARY ===--- Lost some sky time tonight due to (see log for more detail): A large discrepency in mirror numbers for M1's lateral E and F. High wind. New fiducial tables over-wrote existing ones. This forced us to do a pointing model. Engineering camera and guiderMac trouble. However, MARVELS obtained one good exposure on cart 3, plate 3086A and we successfully gathered and installed a new pointing model (in not the best seeing conditions, but the results were ok). ---=== OBSERVING LOG ===--- Afternoon: --------- No spectro focus check made. The spectrographs were last set for an air temp of 3 to 4 C. No data removed from disk. Cartridge 1 mounted. M1 was homed and relaxed to see if actuator B would come back in spec (it was marginally out). It moved from the high side of nominal to the low side. We will continue to monitor it. 00:00Z - Enclosure ventilation started. Night: ----- 00:45Z - Sunset. 00:52Z - Opened enclosure to partly cloudy skies. Force filled spectrographs. 01:13Z - goToField cart 1, MARVELS 3220, field A. Was unable to find an fk5, with several spiral patterns and manual offsets. We checked the mirror numbers, and lateral E and F were off significantly (+2) from the nominal values. We also noticed that the secondary mirror numbers were off as well. We found that collimate was not enabled. Enabling collimate fixed the secondary mirror numbers, but M1 laterals E and F were still off. We homed and relaxed the primary and now everything is happy again. 02:06Z - Evening twilight. 02:46Z - goToField cart1, MARVELS plate 3220, field B. Found fk5 star. Sorry, we forgot to record the first guider frame. 03:11Z - Started 50min MARVELS exposure. Seeing 2.0" 03:15Z - Closed due to high winds. Kept ventilating. 04:45Z - Wind dropped below limit. We mounted cart 3, plate 3086 and then opened the enclosure. 05:03Z - goToField cart 3, MARVELS plate 3086, field A. GSFOGTF. Apologies again, we forgot to record the first guider frame. 05:09Z - 50 min exposure. Seeing 1.7" 06:01Z - goToField for field B. First guider frame gimg0614.fits. We tried everything that we could think of, and we still could not find the guide stars in this field. There were a few random stars coming in and out of the fibers during spiralPatterns, but they definitely were not fk5s. All axes were cleared, fiducials were all ok, mirror numbers were all ok, several offsets and spiral patterns were issued... We decided to try another plate... 06:20Z - Mounted cart 4, MARVELS plate 3219, field B. No guide stars were found on goToField. No fk5 star found either. We were stumped. 06:40Z - Steph called. She was building new new fiducial tables at home when somehow the new ones over-wrote the old ones. We now must obtain a pointing model. We mounted the engineering camera. 08:46Z - After a lot of trouble with setting up the engineering camera and the guiderMac, we finally were able to take an image. The fiber optic cables in the computer room had to be swapped because the paralan recognized A-A and B-B as an error. 08:45Z - Steph rebooted the MCP with the new version of the fiducial tables (v1_79), and reset the PID values. 08:53Z - Homed and relaxed the secondary mirror. Mirror numbers were checked and all were ok. 10:00Z - Began 72-point pointing model. Seeing was not the greatest (over 2", at least... DIMM had trouble even registering Polaris). 11:26Z - Checked and installed the new pointing model. Results below: * getdat pterr_20090210.dat Read 72 data points * fix * fit coeff change value sigma 1 IE +6.36 2 IA -145.16 3 PZZ1 -130.99 4 AN +1.22 5 AW -2.12 6 CA +15.79 7 TX -6.05 8 HWCA2 -0.13 9 HWSA2 -0.25 10 HNCA1 +0.72 11 HNSA1 -1.16 12 ACEC +1.53 13 ACES -1.02 Sky RMS = 20.99 Popn SD = 20.99 * use ie ia pzz1 an aw ca * fit coeff change value sigma 1 IE -22.456 -16.10 0.860 2 IA -3.821 -148.98 1.409 3 PZZ1 -5.007 -136.00 1.059 4 AN +6.110 +7.33 0.339 5 AW -8.203 -10.33 0.339 6 CA -0.473 +15.31 0.983 7 TX -6.05 8 HWCA2 -0.13 9 HWSA2 -0.25 10 HNCA1 +0.72 11 HNSA1 -1.16 12 ACEC +1.53 13 ACES -1.02 Sky RMS = 2.50 Popn SD = 2.61 * use * fit coeff change value sigma 1 IE +0.379 -15.72 1.273 2 IA +0.060 -148.92 1.383 3 PZZ1 +0.832 -135.17 4.154 4 AN -0.191 +7.14 0.434 5 AW +0.096 -10.23 0.425 6 CA -0.228 +15.09 1.024 7 TX -0.400 -6.45 1.978 8 HWCA2 +0.159 +0.03 0.557 9 HWSA2 -0.531 -0.78 0.576 10 HNCA1 -0.407 +0.32 0.649 11 HNSA1 +0.626 -0.54 0.646 12 ACEC +0.289 +1.82 0.783 13 ACES +0.986 -0.03 0.792 Sky RMS = 2.39 Popn SD = 2.64 11:47Z - Began 25-point pointing model to check our results. * getdat pterr_20090210.dat Read 24 data points * use * fit coeff change value sigma 1 IE -2.295 -18.01 3.815 2 IA +4.183 -144.74 10.854 3 PZZ1 -14.841 -150.01 6.670 4 AN -2.399 +4.74 3.957 5 AW +0.490 -9.74 1.586 6 CA -2.114 +12.97 7.944 7 TX +7.229 +0.78 3.122 8 HWCA2 -0.018 +0.01 3.882 9 HWSA2 +1.755 +0.98 2.794 10 HNCA1 -0.801 -0.48 5.466 11 HNSA1 -1.698 -2.23 8.159 12 ACEC +1.648 +3.47 4.667 13 ACES +3.199 +3.16 6.086 Sky RMS = 2.04 Popn SD = 3.02 12:32Z - Morning twilight. 12:35Z - Removed engineering camera and remounted cart 4. Closed enclosure and stopped ventilating. We had to switch back the fiber connections in the computer room in order for the paralan to be happy again. We then rebooted the guiderMac, switched the instrument back to SPECTRO, and then successfully took a guider camera image. 12:46Z - endNight rejoiced. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 7957 23:06Z 23:40Z 100 O -164.18 -155.71 71 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): n QA Procedures Done (y/n): n UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 58781, plate -9999 ------- 00:21 58781 0.0 bias 00:25 58782 0.0 bias ----- sequence 58783, plate 3220 ------- 00:28 58783 0.0 bias ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 05:18Z 3 3086 -81.19 -0.0007 49.00 -0.0032 118.62 0.0260 0.999950 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 05:18Z 3 3086 0.99995 494 1773 180 -81 49.0 -2.9 26 218 - 1.7 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 23:26Z 31 19 22 205 (SSW) 547 - Partly cloudy 23:58Z 29 18 18 194 (SSW) 530 - " 00:29Z 28 17 23 194 (SSW) 413 - " 01:45Z 27 15 22 201 (SSW) 666 - " 02:17Z 28 14 23 197 (SSW) 717 - Clear 02:47Z 28 11 29 198 (SSW) 722 - " 03:17Z 28 10 35 203 (SSW) 700 - " 03:49Z 28 9 27 201 (SSW) 668 - " 04:21Z 28 9 22 218 (SW) 715 - " 04:54Z 27 10 20 209 (SSW) 800 - " 05:25Z 27 11 23 225 (SW) 749 - " 05:55Z 26 9 28 229 (SW) 564 - " 06:26Z 26 7 29 207 (SSW) 541 - " 06:58Z 26 5 26 216 (SW) 473 - " 07:29Z 25 5 22 232 (SW) 433 - " 08:00Z 26 4 27 237 (WSW) 388 - " 08:30Z 25 4 27 226 (SW) 431 - " 09:01Z 25 4 21 225 (SW) 469 - " 09:32Z 25 3 33 233 (SW) 432 - Some light clouds 10:03Z 25 4 23 226 (SW) 441 - Clear 10:34Z 25 4 26 209 (SSW) 414 - " 11:05Z 24 5 23 218 (SW) 675 - " 11:35Z 23 6 22 227 (SW) 699 - " 12:07Z 21 7 18 235 (SW) 559 - " 12:37Z 21 8 18 243 (WSW) 513 - " ---=== TELESCOPE STATUS ===--- 00:00Z - Began ventilation. Doors and louvers opened, fans on. 00:52Z - Opened enclosure to partly cloudy skies. 03:15Z - Closed enclosure, but kept ventilating. Wind gusts 40mph+. 04:45Z - Opened enclosure. Wind fell below limit. 12:35Z - Enclosure is on, doors and louvers are closed, and fans are off. No interlocks bypassed. Status at: 12:46Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 4 Counterweights at: 255 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 144.60 lbs, 15.79 psi Imager 173.30 lbs, 20.54 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_39_0 Watcher: v2_36_0 MCP: v6_4_0 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_43 SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -0.0270 -0.0280 Axial B 0.0530 0.0470 Axial C -0.0330 -0.0360 Trans D -0.7570 -0.7600 Lateral E 10.5918 10.5791 Lateral F 12.1793 12.1793 GALILS Commanded: 10450. -100. 5750. 8600. 10350. 10000. Actual: 10471. -76. 5755. 8616. 10374. 10023. SETMIR VALUES PriDesOrient: 0.00 -11.72 8.60 496.70 640.50 PriOrient: 1.06 -11.65 8.77 495.77 639.59 SECONDARY: ---------- Focus: 0.00 Air Temp.: -1.1 Alt.: 30.363317 MIGS TONIGHT NOMINAL Axial A 18.5830 18.5840 Axial B 15.6190 15.6220 Axial C 13.8150 13.8150 Trans D 15.1070 15.1080 GALILS Commanded: 1626660. 1528821. 1595599. -2950. -4550. Actual: 1626627. 1528808. 1595583. -2945. -4561. SETMIR VALUES SecDesOrient: 1257.00 -12.00 -28.00 -39.00 94.77 SecOrient: 1257.02 -12.00 -28.01 -39.19 94.70 After homing and relaxing M1, PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -0.0310 -0.0280 Axial B 0.0420 0.0470 Axial C -0.0370 -0.0360 Trans D -0.7240 -0.7600 Lateral E 8.7376 10.5791 Lateral F 10.3251 12.1793 GALILS Commanded: 10450. -100. 5750. 8600. -50000. -50000. Actual: 10450. -100. 5750. 8600. -50000. -50000. SETMIR VALUES PriDesOrient: 0.00 -11.72 8.60 496.70 640.50 PriOrient: 1.06 -11.65 8.77 -1334.79 639.59 SECONDARY: ---------- Focus: 0.00 Air Temp.: -2.2 Alt.: 30.352910 MIGS TONIGHT NOMINAL Axial A 18.7200 18.5840 Axial B 15.7520 15.6220 Axial C 13.9430 13.8150 Trans D 15.3830 15.1080 GALILS Commanded: 1458339. 1358440. 1427341. 9100. 7500. Actual: 1458950. 1361150. 1427900. 9100. 7500. SETMIR VALUES SecDesOrient: 1123.75 -12.00 -28.00 -39.00 -182.72 SecOrient: 1122.71 -12.50 -28.55 -39.59 -182.39 After not finding the guide stars, we checked the mirror numbers: PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -0.0300 -0.0280 Axial B 0.0420 0.0470 Axial C -0.0370 -0.0360 Trans D -0.7220 -0.7600 Lateral E 8.7376 10.5791 Lateral F 10.3251 12.1793 GALILS Commanded: 10450. -100. 5750. 8600. -50000. -50000. Actual: 10450. -100. 5750. 8600. -50000. -50000. SETMIR VALUES PriDesOrient: 0.00 -11.72 8.60 496.70 640.50 PriOrient: 1.06 -11.65 8.77 -1334.79 639.59 SECONDARY: ---------- Focus: 0.00 Air Temp.: -2.1 Alt.: 29.999896 MIGS TONIGHT NOMINAL Axial A 18.7200 18.5840 Axial B 15.7540 15.6220 Axial C 13.9470 13.8150 Trans D 15.3850 15.1080 GALILS Commanded: 1453287. 1351857. 1423643. 9100. 7500. Actual: 1458950. 1361150. 1427900. 9100. 7500. SETMIR VALUES SecDesOrient: 1123.75 -12.00 -28.00 -39.00 -182.72 SecOrient: 1118.65 -13.32 -28.86 -39.79 -181.83 We noticed that collimated was not enabled, so we enabled it and checked mirror numbers again: PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -0.0250 -0.0280 Axial B 0.0480 0.0470 Axial C -0.0360 -0.0360 Trans D -0.7290 -0.7600 Lateral E 9.3980 10.5791 Lateral F 11.0744 12.1793 GALILS Commanded: 10450. -100. 5750. 8600. 10350. 10000. Actual: 10450. -100. 5750. 8600. 10350. 10000. SETMIR VALUES PriDesOrient: 0.00 -11.72 8.60 496.70 640.50 PriOrient: 1.06 -11.65 8.77 495.77 639.59 SECONDARY: ---------- Focus: 0.00 Air Temp.: -2.1 Alt.: 30.000008 MIGS TONIGHT NOMINAL Axial A 18.5780 18.5840 Axial B 15.6150 15.6220 Axial C 13.8180 13.8150 Trans D 15.1130 15.1080 GALILS Commanded: 1628002. 1527901. 1593507. -2950. -4550. Actual: 1626650. 1528800. 1595600. -2950. -4550. SETMIR VALUES SecDesOrient: 1257.00 -12.00 -28.00 -39.00 94.77 SecOrient: 1256.57 -11.34 -28.84 -39.74 94.27 Now, the secondary mirror numbers are ok, but M1 lateral E and F are still off. We homed and relaxed the primary one more time, and the results are thus: PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -0.0300 -0.0280 Axial B 0.0260 0.0470 Axial C -0.0370 -0.0360 Trans D -0.7540 -0.7600 Lateral E 10.5664 10.5791 Lateral F 12.1666 12.1793 GALILS Commanded: 10450. -100. 5750. 8600. 10350. 10000. Actual: 10450. -100. 5750. 8600. 10350. 10000. SETMIR VALUES PriDesOrient: 0.00 -11.72 8.60 496.70 640.50 PriOrient: 1.06 -11.65 8.77 495.77 639.59 SECONDARY: ---------- Focus: 0.00 Air Temp.: -1.9 Alt.: 70.000016 MIGS TONIGHT NOMINAL Axial A 18.6020 18.5840 Axial B 15.6340 15.6220 Axial C 13.8500 13.8150 Trans D 15.1110 15.1080 GALILS Commanded: 1598276. 1504745. 1557829. -2950. -4550. Actual: 1626650. 1528800. 1595600. -2950. -4550. SETMIR VALUES SecDesOrient: 1257.00 -12.00 -28.00 -39.00 94.77 SecOrient: 1233.14 -7.76 -27.51 -38.86 91.86 We homed and relaxed the secondary mirror and then checked all of the mirror numbers in preparation for a pointing model: PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -0.0330 -0.0280 Axial B 0.0220 0.0470 Axial C -0.0410 -0.0360 Trans D -0.7570 -0.7600 Lateral E 10.5791 10.5791 Lateral F 12.1793 12.1793 GALILS Commanded: 10450. -100. 5750. 8600. 10350. 10000. Actual: 10450. -100. 5750. 8600. 10350. 10000. SETMIR VALUES PriDesOrient: 0.00 -11.72 8.60 496.70 640.50 PriOrient: 1.06 -11.65 8.77 495.77 639.59 SECONDARY: ---------- Focus: 0.00 Air Temp.: -3.5 Alt.: 29.999997 MIGS TONIGHT NOMINAL Axial A 18.5820 18.5840 Axial B 15.6240 15.6220 Axial C 13.8190 13.8150 Trans D 15.1070 15.1080 GALILS Commanded: 1627876. 1527914. 1595454. -2950. -4550. Actual: 1626650. 1528800. 1595600. -2950. -4550. SETMIR VALUES SecDesOrient: 1257.00 -12.00 -28.00 -39.00 94.77 SecOrient: 1257.06 -12.02 -28.66 -39.62 94.72 ---=== PROBLEMS IN DETAIL ===--- Spectro Focus Check: -------------------- No focus check made. The current actuator values are: sp1> Coll_motor_A -4895 01P-0000004895 Coll_motor_B -5094 02P-0000005094 Coll_motor_C -5014 01P-0000005014 Coll_motor_A_status 133 01S0133 Coll_motor_B_status 133 02S0133 Coll_motor_C_status 133 01S0133 Requested_exp.time 0 Exp_time_left 0 Last_exp.time 2 Last_open_transit.time 700 Last_close_transit.time 525 Humid_Hartmann 12.9 0xFE0F Humid_Cen_Optics 8.7 0xFE46 Temp_Median 5.9 Temp_Hartmann_Top 5.0 (0xFF30) Temp_Red_Cam_Bot 5.7 (0xFF11) Temp_Red_Cam_Top 6.2 (0xFEFD) Temp_Blue_Cam_Bot 6.0 (0xFF03) Temp_Blue_Cam_Top 5.9 (0xFF0A) sp2> Coll_motor_A -4632 01P-0000004632 Coll_motor_B -5181 02P-0000005181 Coll_motor_C -5344 01P-0000005344 Coll_motor_A_status 133 01S0133 Coll_motor_B_status 133 02S0133 Coll_motor_C_status 133 01S0133 Requested_exp.time 0 Exp_time_left 0 Last_exp.time 3 Last_open_transit.time 715 Last_close_transit.time 615 Humid_Hartmann -24.7 0x0 Humid_Cen_Optics 7.6 0xFE54 Temp_Median 3.7 Temp_Hartmann_Top 4.4 (0xFF48) Temp_Red_Cam_Bot 4.1 (0xFF54) Temp_Red_Cam_Top 3.4 (0xFF70) Temp_Blue_Cam_Bot 3.7 (0xFF64) Temp_Blue_Cam_Top 3.9 (0xFF5C) Spectro Bias Check: ------------------ Filename 02% 05% 10% 50% 90% 95% 98% ------------------- ------ ------ ------ ------ ------ ------ ------ sdR-b1-00058780.fit -5.0 -3.9 -2.7 2.6 7.1 8.2 10.4 sdR-b2-00058780.fit -7.9 -6.4 -4.1 0.9 7.2 9.6 11.0 sdR-r1-00058780.fit -6.9 -5.1 -3.6 1.8 6.9 8.3 10.5 sdR-r2-00058780.fit -7.9 -5.8 -4.5 0.7 6.8 8.1 10.2 Done for 71.916346 sec.